OpenStudy (mymomlovesme):

4y-12=-4 Solve for y! I have multiple choice.

OpenStudy (steve816):

Add 12 to both sides first.

alones (alones):

^^4y−12+\color{red}{12}=−4+\color{red}{12} which would be 4y=8 simplified

OpenStudy (mymomlovesme):

wait

OpenStudy (mymomlovesme):

y=11 y=2

alones (alones):

Just like Steve said you add 12 to both sides\[\ 4y−12+\color{red}{12}=−4+\color{red}{12}\] which would be \(4y=8\) simplified

OpenStudy (yournightmare):

y=2

OpenStudy (mymomlovesme):

y=-4 y=-13

OpenStudy (mymomlovesme):

So y=2?

alones (alones):

Yes \(y=2\) would be correct Since you divide \(4\) \[\frac{ 8 }{ 4 }\]

OpenStudy (mymomlovesme):

Okay thanks

alones (alones):

You're welcome cx

OpenStudy (mymomlovesme):

No prob